OpenStudy (anonymous):

Can someone solve this please? the points (2,6) and (3,18) lie on the curve y=ax^n use logarithms to find the values of a and n, giving your answers correct to 2d.p. thanks :)

OpenStudy (anonymous):

6 = a(2^n) ------> a = 6/2^n 18 = a(3^n)-------> 18 = 6(3^n)/(2^n)---->3 = 3^n/2^n

OpenStudy (anonymous):

3 = (3/2)^n

OpenStudy (anonymous):

i dont know what you did :(

OpenStudy (anonymous):

how did you get 3=(3/2)^n?

OpenStudy (anonymous):

because 18 = 6(3/2)^n, divide both sides by 6

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

but i need a number with 2dp :S

OpenStudy (anonymous):

okay. so we have 3 = (3/2)^n take log on both sides log 3 = log(3/2)^n -------> log 3 = n log(3/2) ----> n = log 3/ log(3/2) use calculator to find the value of n to 2 dp

OpenStudy (anonymous):

2.71!!! YAYA! thanks ill give u a medal :)

OpenStudy (anonymous):

you are welcome.

OpenStudy (anonymous):

use n to find a by substituting in one of the equations.

OpenStudy (anonymous):

oh :( i thought that the qs ended ill do that now then :)

myininaya (myininaya):

ok I got n=log(1/3) /log(2/3) maybe I made a mistake

OpenStudy (anonymous):

:S

myininaya (myininaya):

oh wait it might be the same thing lol

OpenStudy (anonymous):

ya lol

myininaya (myininaya):

there are so many properties with log

myininaya (myininaya):

the reason they are the same is : log(1/3) /log(2/3)=[log1-log3]/[log2-log3]=[0-log3]/[(-1)(log3-log2)]=log3/[log3/2]

OpenStudy (anonymous):

wow, it looks like a sentence without any words XD

OpenStudy (anonymous):

its not hard to understand if you think about it, Miss.

OpenStudy (anonymous):

n=ln3\(ln3/2) et a=6/[2^(ln3/2)]

OpenStudy (anonymous):

i dont actually understand it it looks like random equations put together :p

OpenStudy (anonymous):

okay. I hope for your sake that you are not taking math as a major course then :)

OpenStudy (anonymous):

no, really wanna drop it next year. its way too hard i cant manage it :(

OpenStudy (anonymous):

yes, some people are good at certain things, others are good at other things.